You are pulling a 10 kg object with a 87 N force w
Answer #1

here n = mg-fsin60 = 22.66 N

w1 = 87*0.51*cos60 = 22.19 J

w2 = -0.5*22.66*0.51 = -5.78 J

w tot = w1+w2 = 22.19-5.78

wtot = 16.41 J